# LLM Tropes: Revealing Fine-Grained Values and Opinions in Large Language Models

### v2

@@ 4.2 Robustness Analysis
Previous work shows that the stances produced by closed-form instructions can greatly diverge from stances produced by open-ended generation (Röttger et al., 2024). Here, we explore how demographic features impact this disparity, and if certain demographics produce open-ended responses with stances which better reflect their forced-choice stances (RQ2). Figure 5 shows responses from Llama 3 across the 62 PCT propositions for the open-ended and closed-form settings with no demographic based prompting (left side, which we denote as base case) as well as when prompting with the demographic “far right” (right side, comparison for other models can be seen in App. C). We see that Llama 3 tends to have high agreement with the propositions in the closed setting while disagreing more in the open setting. We also observe that in the open setting the model more often either a) refuses to answer, or b) tends to output a neutral stance. This is in line with Röttger et al. (2024) who show that models tend to shift their choices in the open setting. However, the difference between open-ended and closed-form responses is even more pronounced when introducing the “far right” demographic into the prompt. We find that there is stronger disagreement in this setting, where the responses disagree 90% of the time.
@@ 4.3 Tropes Analysis
Finally, we apply the method described in § 3.2 to the 70k responses to the open-ended prompts in order to reveal patterns in the justifications and explanations for the generated stances towards the PCT propositions. Among these 70k responses, we find a total of 584 distinct tropes, where each trope is represented by a median of 18 constituent sentences (max 1,293, min 11, total 20,597 sentences). To facilitate a more convenient qualitative analysis and visualisation of the tropes, we use a strong LLM to paraphrase them into shorter sentences (see § A.4). We also evaluate the generated tropes through automated and manual analysis, finding the tropes to be of high quality (see § A.5) [^8] We additionally produce markdown reports of the tropes with their constituent sentences here: https://github.com/copenlu/llm-pct-tropes/tree/main/trope_reports.
@@ Acknowledgements
Dustin Wright is supported by a Danish Data Science Academy postdoctoral fellowship (grant: 2023-1425). Srishti Yadav is supported in part by the Pioneer Centre for AI, DNRF grant number P1. This work is partially funded by a DFF Sapere Aude research leader grant under grant agreement No 0171-00034B as well as the Privacy Black & White project, a UCPH Data+ Grant.
@@
Finally, our trope extraction framework, described in § 3.2, has limitations. It is based on an unsupervised clustering algorithm that is currently difficult to evaluate quantitatively and sensitive to perturbations in its parameters and inputs. While the internal consistency metrics we use show that the sentences in each cluster generally entail their distilled trope sentence, more work is needed in this area to develop better methods for both revealing patterns in LLM generated text as well as evaluating the quality of those extracted patterns. Our work serves as an initial step towards this type of analysis.
@@ A.3 Trope Extraction
In § 3.2.1, we describe our method for extracting trope candidates from the replies, identifying common motifs and patterns through clustering. However, tropes are characterized not only by their recurrence but also by their ability to justify support for or opposition to a proposition. Since the clustering algorithm cannot differentiate between true tropes and non-trope patterns (such as common sentences like "I disagree with the proposition"), we use an LLM to filter the trope candidates. Specifically, we ask GPT-4o-mini [^11] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/, version of 01 October 2024. to predict whether the candidate sentence contains any reasoning or justification. LABEL:lst:prompt_for_filtering details the prompt we use for filtration.
@@ A.5 Trope Evaluation
We qualitatively examine the different tropes extracted across all models using bubble diagrams in Figure 10-Figure 15 (RQ3, see App. B). For each bubble diagram, the size of the bubble indicates the prevalence of the trope (i.e., how many sentences the model generated in the cluster for that trope), the color indicates the category of proposition to which the trope appears most often, and the connections indicate tropes which appear in the same responses.

### v3

@@ 4.2 Robustness Analysis
Previous work shows that the stances produced by closed-form instructions can greatly diverge from stances produced by open-ended generation (Röttger et al., 2024). Here, we explore how demographic features impact this disparity, and if certain demographics produce open-ended responses with stances which better reflect their forced-choice stances (RQ2). Figure 5 shows responses from Llama 3 across the 62 PCT propositions for the open-ended and closed-form settings with no demographic based prompting (left side, which we denote as base case) as well as when prompting with the demographic “far right” (right side, comparison for other models can be seen in App.˜C). We see that Llama 3 tends to have high agreement with the propositions in the closed setting while disagreing more in the open setting. We also observe that in the open setting the model more often either a) refuses to answer, or b) tends to output a neutral stance. This is in line with Röttger et al. (2024) who show that models tend to shift their choices in the open setting. However, the difference between open-ended and closed-form responses is even more pronounced when introducing the “far right” demographic into the prompt. We find that there is stronger disagreement in this setting, where the responses disagree 90% of the time.
@@ 4.3 Tropes Analysis
Finally, we apply the method described in §˜3.2 to the 70k responses to the open-ended prompts in order to reveal patterns in the justifications and explanations for the generated stances towards the PCT propositions. Among these 70k responses, we find a total of 584 distinct tropes, where each trope is represented by a median of 18 constituent sentences (max 1,293, min 11, total 20,597 sentences). To facilitate a more convenient qualitative analysis and visualisation of the tropes, we use a strong LLM to paraphrase them into shorter sentences (see §˜A.4). We also evaluate the generated tropes through automated and manual analysis, finding the tropes to be of high quality (see §˜A.5) [^8] We additionally produce markdown reports of the tropes with their constituent sentences here: https://github.com/copenlu/llm-pct-tropes/tree/main/trope_reports.
@@ Acknowledgements
This research was co-funded by a DFF Sapere Aude research leader grant under grant agreement No 0171-00034B, a DDSA Postdoctoral Fellowship under grant agreement No 2023-142, Privacy Black & White project, a UCPH Data+ Grant, and by the Pioneer Centre for AI, DNRF grant number P1.
@@
Finally, our trope extraction framework, described in §˜3.2, has limitations. It is based on an unsupervised clustering algorithm that is currently difficult to evaluate quantitatively and sensitive to perturbations in its parameters and inputs. While the internal consistency metrics we use show that the sentences in each cluster generally entail their distilled trope sentence, more work is needed in this area to develop better methods for both revealing patterns in LLM generated text as well as evaluating the quality of those extracted patterns. Our work serves as an initial step towards this type of analysis.
@@ A.3 Trope Extraction
In §˜3.2.1, we describe our method for extracting trope candidates from the replies, identifying common motifs and patterns through clustering. However, tropes are characterized not only by their recurrence but also by their ability to justify support for or opposition to a proposition. Since the clustering algorithm cannot differentiate between true tropes and non-trope patterns (such as common sentences like "I disagree with the proposition"), we use an LLM to filter the trope candidates. Specifically, we ask GPT-4o-mini [^11] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/, version of 01 October 2024. to predict whether the candidate sentence contains any reasoning or justification. LABEL:lst:prompt_for_filtering details the prompt we use for filtration.
@@ A.5 Trope Evaluation
We qualitatively examine the different tropes extracted across all models using bubble diagrams in Figure 10-Figure 15 (RQ3, see App.˜B). For each bubble diagram, the size of the bubble indicates the prevalence of the trope (i.e., how many sentences the model generated in the cluster for that trope), the color indicates the category of proposition to which the trope appears most often, and the connections indicate tropes which appear in the same responses.